# A Survey of Generative AI for Intelligent Transportation Systems

### v1

@@ Abstract.
Intelligent transportation systems play a crucial role in modern traffic management and optimization, greatly improving traffic efficiency and safety. With the rapid development of generative artificial intelligence (Generative AI) technologies in the fields of image generation and natural language processing, generative AI has also played a crucial role in addressing key issues in intelligent transportation systems, such as data sparsity, difficulty in observing abnormal scenarios, and in modeling data uncertainty. In this review, we systematically investigate the relevant literature on generative AI techniques in addressing key issues in different types of tasks in intelligent transportation systems. First, we introduce the principles of different generative AI techniques, and their potential applications. Then, we classify tasks in intelligent transportation systems into four types: traffic perception, traffic prediction, traffic simulation, and traffic decision-making. We systematically illustrate how generative AI techniques addresses key issues in these four different types of tasks. Finally, we summarize the challenges faced in applying generative AI to intelligent transportation systems, and discuss future research directions based on different application scenarios.
@@ 1. Introduction
The transportation system contains a range of traffic infrastructure, equipment, and management strategies, including roads, traffic control, and traffic planning, designed to manage traffic flow and meet people’s travel needs. Its efficiency significantly impacts both societal and the economic activities. The rapid urbanization has led to a steep increase in the number of vehicles, resulting in prevalent issues like congestion and accidents. These challenges inconvenience travelers and disrupt urban operations, emphasizing the urgency of traffic management and optimization. In recent years, the rapid advancement of computer technology has given rise to Intelligent Transportation Systems (ITS). ITS harnesses cutting-edge technologies, such as artificial intelligence, big data analysis, and the Internet of Things, to create a comprehensive system integrating people, roads, and vehicles. Its goals include reducing congestion, enhancing safety, conserving energy, lowering carbon emissions, and supporting autonomous driving technology.
A fundamental component for realizing ITS is the effective utilization of traffic data. For traffic authorities, they benefit from real-time data collected through sensors, cameras, and various devices, enabling them to monitor and manage road conditions. Furthermore, analyzing historical traffic data can be helpful for predictions regarding future traffic flow and congestion. For travelers, they benefit from real-time traffic data through navigation apps and traffic management systems, which provide up-to-date traffic information to assist in route selection and congestion avoidance. In addition, autonomous vehicles heavily rely on numerous sensors and data to perceive their surroundings, make effective decisions for safe and efficient driving. These aspects indicate the important of traffic data within intelligent transportation systems.
@@
In this context, traditional deep learning methods have been widely adopted due to their efficient application in traffic data analysis and modeling. These algorithms can process massive amounts of traffic data, automatically recognize patterns, predict trends, and optimize decisions, thereby improving traffic efficiency and safety. For example, traditional deep learning techniques such as convolutional neural networks (CNNs) are used to analyze traffic camera images, further identifying vehicles, pedestrians, traffic signs, and road conditions to monitor traffic flow and situations. Recurrent neural networks (RNNs) are suitable for processing sequence data, such as time series data generated by traffic sensors, and can be used to predict traffic flow, congestion, and the likelihood of accidents. Deep reinforcement learning (DRL) methods are applied to the development of autonomous driving technology, training models to make driving decisions by simulating the behavior of intelligent agents (such as autonomous vehicles) in traffic environments. These models can enable safe autonomous driving. By utilizing these deep learning models, intelligent transportation systems can analyze road condition patterns more effectively, predict potential congestion points, and provide adaptive decision support for optimizing traffic efficiency.
@@
However, with the increasing complexity of transportation networks, the interweaving of various modes of transportation, and the growing demand for travel, the traffic environment has become increasingly intricate. This complexity poses numerous challenges for traditional deep learning methods in addressing traffic-related issues.
@@
Generative Artificial Intelligence (Generative AI) technology has recently developed rapidly. It can automatically generate various forms of content based on user input, demands, or instructions, including but not limited to text, images, videos, audio, etc. The emergence of generative AI technology makes content production more efficient and convenient, and has broad application prospects. In the field of generative AI, many mature technologies have been widely applied. For example, Variational Auto-Encoder (VAE) (Kingma and Welling, 2013) is a generative adversarial network that can model and generate complex data. Generative Adversarial Network (GAN) (Goodfellow et al., 2020) improves the quality of generated content through the adversarial process of two neural networks (i.e., generators and discriminators). Normalizing Flow (Rezende and Mohamed, 2015) is a normalization technique used for generating complex data distributions, while Energy-Based Model (LeCun et al., 2006) is a generative model based on energy that can effectively improve the quality of generation. Additionally, Generative Model from Physical Process (Liu et al., 2023c) is a generative model based on physical processes that can be used to generate content with physical constraints. Diffusion Model (Sohl-Dickstein et al., 2015) is a probabilistic method that generates data by adding noise to data gradually to reduce its quality. Generative Pre-trained Transformer (GPT) (Ghojogh and Ghodsi, 2020) is based on the Transformer architecture (Vaswani et al., 2017), which pre-trains a large amount of text data and then uses it to generate new text content.
These generative AI techniques offer significant advantages in data generation, transformation, and repair, and they can effectively model the uncertainty of data. As a result, these technologies have garnered significant attention in different fields. Especially in the field of intelligent transportation, generative AI technology has played a crucial role and introduced innovative solutions to address traffic-related challenges (Lv, 2023). Numerous researchers are actively engaged in the application of diverse generative AI techniques to traffic-related tasks. For example, in autonomous driving scenarios, generative AI technology proves invaluable in generating high-fidelity driving scene images and videos, which are essential for the training and testing of autonomous driving systems (Krajewski et al., 2018; Kumar et al., 2023). Utilizing generative AI-generated simulation scenarios, autonomous vehicles can conduct extensive driving simulations within virtual environments, thereby enhancing their ability to make precise real-world decisions. In traffic flow prediction tasks, generative AI technology can learn and model the distribution of traffic data, thus generating future traffic flow data (Zhang et al., 2022b; Zand et al., 2023; Wen et al., 2023). This is very helpful for traffic authorities to carry out effective traffic planning and optimization.
@@
Blockchain (Swain, 2006; Das et al., 2023), Edge computing (Zhou et al., 2021a; Gong et al., 2023), Dynamic pricing (Saharan et al., 2020), Fuzzy logic (Swain, 2006),
Swarm intelligence (Mathew, 2020), V2X downloading (Wang et al., 2019), Data mining (Anand et al., 2018; El Faouzi et al., 2011; Zhu et al., 2018),
@@
Driver behavior detection (Chhabra et al., 2017), Vehicle classification (Gholamhosseinian and Seitz, 2021; Yang and Pun-Cheng, 2018), Urban mobility (Mangiaracina et al., 2017),
Urban monitoring (Lima et al., 2017), Traffic sign detection (Ellahyani et al., 2021; Luo et al., 2020), Traffic prediction (Suhas et al., 2017; Yuan and Li, 2021),
Autonomous driving (Altaf and Kaul, 2021), Metaverse (Njoku et al., 2023), Trust management (Ma et al., 2011), Spectrum regulation (Choi et al., 2020)
@@
3D (Li et al., 2023d), Biology (Zhang et al., 2023b), ChatGPT (Zhang et al., 2023e; Wang et al., 2023a), Metaverse (Qin and Hui, 2023), Edge cloud computing (Wang et al., 2023b),
Mobile network (Xu et al., 2023a; Du et al., 2023), Security & Privacy (Wang et al., 2023c; Chen et al., 2023b), Vehicle network (Zhang et al., 2023d)
@@
Table 1. Related reviews of intelligent transportation system (ITS) and generative AI. DRL: Deep reinforcement learning, GNN: Graph neural network, NLP: Natural language processing, V2X: Vehicle to everything. DL: Deep learning. IoT: Internet of things. VANET: Vehicle ad-hoc network.
As shown in Table 1, there are numerous surveys on intelligent transportation systems or generative AI. However, most of them rarely explore the practical applications of generative AI technology in intelligent transportation systems. Although the authors in (Lin et al., 2023a) have explored the role of GANs in intelligent transportation systems, they do not comprehensively cover other generative AI techniques. Importantly, they do not systematically examine the advantages of generative AI techniques over traditional deep learning methods in intelligent transportation systems. Therefore, in this paper, we will explore the applications of generative AI techniques in traffic perception, prediction, simulation, and decision-making within intelligent transportation systems, as shown in Figure 1. Our goal is to provide a comprehensive analysis of their role in promoting intelligent transportation, and offer a constructive insight to help readers gain a deeper understanding of the challenges and opportunities of generative AI in intelligent transportation systems. We expect this survey to provide useful insights for further research and advancement of intelligent transportation systems.
@@
The structure of this paper is as follows. Section 2 introduces mainstream generative AI techniques and their application achievements in different fields, followed by a detailed comparative analysis. The following four sections respectively discuss the applications of generative AI technology in traffic perception, traffic prediction, traffic simulation, and traffic decision-making. Finally, in Section 7, we introduce the challenges faced by generative AI in the application of intelligent transportation systems, and look forward to future research directions.
2. Technologies, Applications and Advantages of Generative AI
As described in previous works (Cao et al., 2023; Foo et al., 2023; Wang et al., 2023a; Wu et al., 2023), generative AI refers to a subset of AI algorithms that focus on creating or producing personalized, and high-quality content, including text, images, videos, and 3D assets, in response to user input or specific requirements. This section provides a comprehensive overview of the underlying technologies, highlights their significant applications, and summarizes the advantages of generative AI techniques in comparison to traditional deep learning methods.
@@ 2.1. Key Technologies
With the rapid growth of computer technology and computing power, generative AI technology has witnessed significant progress over the past decade. Variational Autoencoders (VAE), a probabilistic graphical model, was pioneered by Kingma and Welling in 2013 for data generation (Kingma and Welling, 2013). In 2014, Ian Goodfellow et al. designed generative adversarial networks (GANs) (Goodfellow et al., 2014), which led to a major transformation in the field of generative AI. In 2018, OpenAI proposed Generative Pre-trained Transformer (GPT), which also brought important breakthroughs to the field. Beyond these approaches, technologies like Normalizing Flow, Energy-Based Models (EBMs), Diffusion Probabilistic Models (DPMs) and Generative Models from Physical Process (GenPhys) have also demonstrated substantial potential in the domains of text and image generation. Figure 2 depicts the evolution timeline of generative AI techniques, including VAE, GANs, Normalizing Flow, EBMs, DPMs, GenPhys, and GPT. In this subsection, we will provide an initial introduction to these techniques, followed by a comparative analysis.
@@ 2.1.1. VAE
VAE (Kingma and Welling, 2013) aims to learn the latent representations of data and generate new data samples. It achieves this by mapping data to a probability distribution in a latent space and then re-mapping the latent variables back to the data space through a decoder.
@@ 2.1.2. GAN
GAN (Goodfellow et al., 2014) consists of a generator and a discriminator. They are trained together through an adversarial process to generate realistic data samples. The generator $G$ takes random noise as input, drawn from a prior distribution $p_{z}(z)$, and generates fake data samples $G(z)$ that resemble real data samples. The discriminator $D$ receives both real data samples $x$ from the dataset following $p_{\text{data }}(\boldsymbol{x})$ and fake data samples $G(z)$ generated by the generator. Its objective is to differentiate between real and fake samples. The objective of GAN is to find a Nash equilibrium where the generator generates data that is indistinguishable from real data. This equilibrium is achieved when the discriminator cannot reliably distinguish between real and fake samples.
@@ 2.1.3. Normalizing Flows
The core idea of normalizing flows (Rezende and Mohamed, 2015) is to map a simple probability distribution to a complex one by applying a sequence of reversible transformations, so as to generate samples that match the complex data distribution. These reversible transformations can be combined using the chain rule to efficiently compute the logarithmic probability density of the generated samples. The key feature is that these transformations are all reversible, allowing for easy mapping from the data space to the latent space and vice versa.
Let $f$ be an invertible and smooth mapping function, and its inverse function $f^{-1}=g$. By transforming a random variable $\mathbf{z}\sim p(\mathbf{z})$ using $f$, the distribution of the resulting random variable $\mathbf{y}=f(\mathbf{z})$ is:
(1)1( 1 ) $p\left(\mathbf{y}\right)=p(\mathbf{z})\left|\operatorname{det}\frac{\partial f^{-1}}{\partial\mathbf{y}}\right|=p(\mathbf{z})\left|\operatorname{det}\frac{\partial f}{\partial\mathbf{z}}\right|^{-1},$
where the chain rule is applied. By combining multiple simple mappings and successively using Equation 1, arbitrarily complex densities can be derived. The density $p_{K}(\mathbf{z})$ calculated by the successive transformations of a random variable $\mathbf{z}_{0}$ following the distribution $p_{0}$ through a chain of $K$ transformations is written as:
(2)2( 2 ) $\mathbf{z}_{K}=f_{K}\circ\ldots\circ f_{2}\circ f_{1}\left(\mathbf{z}_{0}\right),$
(3)3( 3 ) $\ln p_{K}\left(\mathbf{z}_{K}\right)=\ln p_{0}\left(\mathbf{z}_{0}\right)-\sum_{k=1}^{K}\ln\left|\operatorname{det}\frac{\partial f_{k}}{\partial\mathbf{z}_{k-1}}\right|.$
Here, the path that is traversed by the random variables $\mathbf{z}_{k}=f_{k}(\mathbf{z}_{k-1})$, which are generated from the initial distribution $p_{0}(\mathbf{z}_{0})$, is known as the flow. The path created by the consecutive distributions $p_{k}$ is referred to as a normalizing flow. By selecting suitable transformations $f_{k}$, we can begin with simple, factored distributions like an independent Gaussian. Then, by applying normalizing flows of varying lengths, we can obtain increasingly complex and multi-modal distributions.

### v2

@@ Abstract.
Intelligent transportation systems are vital for modern traffic management and optimization, greatly improving traffic efficiency and safety. With the rapid development of generative artificial intelligence (Generative AI) technologies in areas like image generation and natural language processing, generative AI has also played a crucial role in addressing key issues in intelligent transportation systems (ITS), such as data sparsity, difficulty in observing abnormal scenarios, and in modeling data uncertainty. In this review, we systematically investigate the relevant literature on generative AI techniques in addressing key issues in different types of tasks in ITS tailored specifically for road transportation. First, we introduce the principles of different generative AI techniques. Then, we classify tasks in ITS into four types: traffic perception, traffic prediction, traffic simulation, and traffic decision-making. We systematically illustrate how generative AI techniques addresses key issues in these four different types of tasks. Finally, we summarize the challenges faced in applying generative AI to intelligent transportation systems, and discuss future research directions based on different application scenarios.
@@ 1. Introduction
The transportation system contains traffic infrastructure, equipment, and management strategies, including traffic control and traffic planning, all aimed at managing traffic flow and meet travel demands. Its efficiency significantly affects societal and economic activities. Rapid urbanization has greatly increased the number of vehicles, leading to congestion and accidents. These challenges inconvenience travelers and disrupt urban operations, emphasizing the need of better traffic management. Recent advances in computing technology have enabled the development of Intelligent Transportation Systems (ITS). Traditionally, ITS covers air, rail, road, and marine transportation. In this paper, we focus on urban road transportation and define ITS as an integrated system connecting people, roads, and vehicles.
The foundational framework of ITS consists of four main components, including traffic perception, prediction, simulation, and decision-making. Traffic perception collects and analyzes large-scale data through sensors, cameras, and other monitoring devices to understand traffic dynamics, which in turn supports prediction models for forecasting future demands and conditions. Traffic simulation models traffic patterns under different scenarios, concentrating on realistic traffic state distributions to validate decision-making feasibility. The decision-making component leverage insights from perception, prediction, and simulation to implement effective traffic control measures and optimize transportation operations, while also generating new traffic data to evaluate their impact. Thus, these components form a cohesive closed-loop system that addresses the majority of ITS applications, enabling effective management, control, and optimization of transportation networks (Lin et al., 2023a; Zhao et al., 2023).
@@
Traditional deep learning methods have been widely adopted in ITS. These methods can process large number of traffic data, automatically recognize patterns, predict trends, and optimize decisions, thereby improving traffic efficiency and safety. For example, convolutional neural networks (CNNs) are used to analyze traffic camera images, identifying vehicles, pedestrians, traffic signs, and road conditions to monitor traffic flow and situations. Recurrent neural networks (RNNs) are effective for processing sequential data, such as time series data from traffic sensors, enabling predictions of traffic flow, congestion, and accident risk. Deep reinforcement learning (DRL) is applied to the field of autonomous driving, where models are trained to make effective driving decisions by simulating the behavior of intelligent agents, such as autonomous vehicles, in traffic environments. By utilizing these deep learning techniques, ITS can more effectively analyze road conditions, predict congestion, and provide adaptive decision support to optimize traffic efficiency.
@@
However, with the increasing complexity of transportation networks, the integration of multiple transportation modes, and the growing demand for travel have made the traffic environment more complex. This complexity poses numerous challenges for traditional deep learning methods in addressing traffic-related issues.
@@
Generative Artificial Intelligence (Generative AI) technology has recently developed rapidly in recent years, enabling the automatic generation of content, including text, images, videos, and audio, based on user input or instructions. Several established techniques play key roles in generative AI. For example, Variational Auto-Encoder (VAE) (Kingma and Welling, 2013) is a generative adversarial network that models and generates complex data. Generative Adversarial Network (GAN) (Goodfellow et al., 2020) improves the quality of generated content through the adversarial process of generators and discriminators. Other techniques include normalizing flow (Rezende and Mohamed, 2015) for generating complex data distributions, energy-based models (LeCun et al., 2006) for enhancing generation quality, and generative models from physical processes (Liu et al., 2023c) for generating content with physical constraints. Additionally, diffusion probabilistic models (Sohl-Dickstein et al., 2015) generate data by adding noise to data gradually to reduce its quality. Generative Pre-trained Transformer (GPT) (Ghojogh and Ghodsi, 2020) use pre-trained Transformer architectures(Vaswani et al., 2017) to generate text.
These generative AI techniques offer significant advantages in data generation, reconstruction, and data uncertainty modeling, which hold potential for addressing challenges in ITS (Lv, 2023). For example, in autonomous driving scenarios, generative AI can generate high-fidelity driving scene images and videos, which are crucial for training and testing autonomous systems (Krajewski et al., 2018; Kumar et al., 2023). By using AI-generated simulation scenarios, autonomous vehicles can perform extensive driving simulations in virtual environments, improving their ability to make accurate real-world decisions. In traffic flow prediction tasks, generative AI technology can learn and model the distribution of traffic data, thus generating future traffic flow data (Zhang et al., 2022b; Zand et al., 2023; Wen et al., 2023). This capability helps traffic authorities improve traffic planning and optimization.
@@
Diffusion models (Peng et al., 2024), Blockchain (Swain, 2006; Das et al., 2023), Edge computing (Zhou et al., 2021a; Gong et al., 2023), Dynamic pricing (Saharan et al., 2020),
Fuzzy logic (Swain, 2006), Swarm intelligence (Mathew, 2020), V2X downloading (Wang et al., 2019), Data mining (Anand et al., 2018; El Faouzi et al., 2011; Zhu et al., 2018),
@@
Driver behavior detection (Chhabra et al., 2017), Vehicle classification (Gholamhosseinian and Seitz, 2021; Yang and Pun-Cheng, 2018), Metaverse (Njoku et al., 2023), Urban mobility (Mangiaracina et al., 2017),
Urban monitoring (Lima et al., 2017), Traffic sign detection (Ellahyani et al., 2021; Luo et al., 2020), Traffic prediction (Suhas et al., 2017; Yuan and Li, 2021), Traffic
simulation (Chen et al., 2024b), Autonomous driving (Altaf and Kaul, 2021; Cui et al., 2024c), Trust management (Ma et al., 2011), Spectrum regulation (Choi et al., 2020)
@@
3D (Li et al., 2023e), Biology (Zhang et al., 2023b), ChatGPT (Zhang et al., 2023e; Wang et al., 2023a), Metaverse (Qin and Hui, 2023), Edge cloud computing (Wang et al., 2023b), Mobile
network (Xu et al., 2023a; Du et al., 2023), Security & Privacy (Wang et al., 2023c; Chen et al., 2023b), Vehicle network (Zhang et al., 2023d)
@@
Table 1. Related reviews of ITS and generative AI. DRL: Deep reinforcement learning, GNN: Graph neural network, NLP: Natural language processing, V2X: Vehicle to everything. DL: Deep learning. IoT: Internet of things. VANET: Vehicle ad-hoc network.
Given the growing potential of generative AI in ITS, it is crucial to conduct an extensive review of past research that focuses on generative AI techniques applied to ITS. As shown in Table 1, numerous surveys have been conducted on ITS or generative AI. However, most of these studies do not delve deeply into the practical applications of generative AI in ITS. For example, while one study focuses on GANs (Lin et al., 2023a) and another on diffusion models (Peng et al., 2024) in ITS, neither comprehensively addresses other generative AI techniques. Moreover, these surveys do not systematically examine how generative AI can overcome challenges faced by traditional deep learning approaches in ITS. To bridge this gap, this paper explores the applications of generative AI techniques in traffic perception, prediction, simulation, and decision-making within ITS, as shown in Figure 1. Our goal is to provide a comprehensive analysis of their role in promoting intelligent transportation, and offer a constructive insight to help readers gain a deeper understanding of the challenges and opportunities of generative AI in intelligent transportation systems. We expect this survey to provide useful insights for further research and advancement of intelligent transportation systems.
Our survey focused on high-quality literature from reputable conferences, books, and journals. We referred to established archives like arXiv and Google Scholar, and major databases such as IEEE Xplore, ACM Digital Library, Springer, and ScienceDirect. Keywords like ”Intelligent Transportation System” combined with generative AI terms (e.g., ”variational autoencoder,” ”generative adversarial network,” ”GPT,” ”LLM”) were used to identify relevant references. Additional filters included terms like ”traffic signal,” ”traffic prediction,” ”traffic simulation,” and ”data imputation.” We screened literature by examining titles, eliminating low-quality papers based on citation count, author affiliation, and publication authority, followed by an abstract review to assess contributions. This process ensured a foundation of credible, high-quality research.
@@
The structure of this paper is as follows. Section 2 introduces mainstream generative AI techniques and gives a detailed comparative analysis. The following four sections respectively discuss the applications of generative AI technology in traffic perception, traffic prediction, traffic simulation, and traffic decision-making. Finally, in Section 7, we introduce the challenges faced by generative AI in the application of ITS, and look forward to future research directions.
2. Fundamental Generative AI Technologies
As described in previous works (Cao et al., 2023; Foo et al., 2023; Wang et al., 2023a; Wu et al., 2023), generative AI refers to a subset of AI algorithms that create personalized and high-quality content such as text, images, videos, and 3D assets based on user input or specific requirements. This section provides an overview of the fundamental technologies behind generative AI.
@@ 2.1. Key Technologies
With the rapid advancement of computer technology and computing power, generative AI has made significant progress over the past decade. In 2013, Kingma and Welling introduced VAE (Kingma and Welling, 2013), a probabilistic graphical model designed for data generation. In 2014, Ian Goodfellow et al. designed GANs (Goodfellow et al., 2014), which had a profound impact on the field of generative AI. In 2018, OpenAI’s GPT made significant contributions to the field. Beyond these, technologies like Normalizing Flow, Energy-Based Models (EBMs), Diffusion Probabilistic Models (DPMs) and Generative Models from Physical Process (GenPhys) have shown substantial potential for text and image generation. Figure 2 depicts the evolution timeline of generative AI techniques, including VAE, GANs, Normalizing Flow, EBMs, DPMs, GenPhys, and GPT. In this subsection, we will provide an initial introduction to these techniques, followed by a comparative analysis.
@@ 2.1.1. VAE
Variational Autoencoders (VAEs) (Kingma and Welling, 2013) aim to learn latent representations of data and generate new samples by mapping data to a latent space and using a decoder to reconstruct it.
@@
In ITS applications, this approach allows capturing latent patterns in traffic data, generating synthetic traffic data that can be used for simulation, traffic prediction, and traffic decision-making.
@@ 2.1.2. GAN
Generative Adversarial Networks (GANs) (Goodfellow et al., 2014) consist of a generator and a discriminator, trained together in an adversarial manner to create realistic data samples. The generator creates fake data from random noise, while the discriminator tries to distinguish it from real data The goal is to reach a point where the generator produces data that the discriminator cannot reliably differentiate from real data, achieving a Nash equilibrium. This adversarial process enables ITS researchers to generate high-quality synthetic traffic scenarios that can be used for various applications, e.g., vehicle control. Generative Adversarial Imitation Learning (GAIL) extends GANs to learn policies by imitating expert behaviors without requiring explicit reward signals (Ho and Ermon, 2016).
@@ 2.1.3. Normalizing Flows
Normalizing flows (Rezende and Mohamed, 2015) transform a simple probability distribution into a complex one by applying a series of invertible transformations. These transformations are reversible, allowing for efficient calculation of probability densities.
Let $f$ be an invertible function that maps a random variable $\mathbf{z}$, sampled from a simple distribution $p(\mathbf{z})$, to a new variable $\mathbf{y}=f(\mathbf{z})$. The probability density of $\mathbf{y}$ is given by: $p\left(\mathbf{y}\right)=p(\mathbf{z})\left|\operatorname{det}\frac{\partial f^{-1}}{\partial\mathbf{y}}\right|=p(\mathbf{z})\left|\operatorname{det}\frac{\partial f}{\partial\mathbf{z}}\right|^{-1}.$ By applying $K$ transformations to an initial variable $\mathbf{z}_{0}\sim p_{0}(\mathbf{z}_{0})$, we obtain a final variable $\mathbf{z}_{K}$ and its log probability density: $\ln p_{K}\left(\mathbf{z}_{K}\right)=\ln p_{0}\left(\mathbf{z}_{0}\right)-\sum_{k=1}^{K}\ln\left|\operatorname{det}\frac{\partial f_{k}}{\partial\mathbf{z}_{k-1}}\right|.$
This allows starting from a simple distribution, like a Gaussian, and creating more complex, multi-modal distributions. In the context of ITS, this allows modeling complex traffic data distributions while maintaining tractable probability densities. By applying normalizing flows to a simple distribution, we can accurately estimate the likelihood of specific traffic conditions, which is crucial for tasks like anomaly detection and probabilistic traffic forecasting.